OpenStudy (anonymous):

Quadratic formula: 4x^2 + 9 = 12x How do I solve this equation, if there is no "c" -b sqroot b^2-4ac / 2a

OpenStudy (anonymous):

\[4x^2 + 9 = 12x \\ 4x^2-12x+9=0\]

OpenStudy (anonymous):

quadratic formula is for the form \[ax^2+bx+c=0\] only

OpenStudy (anonymous):

So can you elaborate how exactly you did that and why?

OpenStudy (anonymous):

Oh wait so you change it into that equation/form

OpenStudy (anonymous):

the "why" is a good question, which i hope i answered with the second post

OpenStudy (anonymous):

the "how" is starting with \[4x^2 + 9 = 12x\] i subtracted \(12x\) from both sides, to put it in the necessary form

OpenStudy (anonymous):

I got a -37.5 Which I'm thinking is wrong?

OpenStudy (anonymous):

+4 -28 ?